# Positive formula for the product of conjugacy classes on the unitary group

### v1

@@ 3.1. A first density formula
$\mathrm{d}\mathbb{P}[\gamma|\alpha,\beta]=$ $\frac{\Delta(\mathrm{e}^{2i\pi\gamma})sf(n-1)}{(2\pi)^{n-1}n!\Delta(\mathrm{e}^{2i\pi\alpha})\Delta(\mathrm{e}^{2i\pi\beta})}\sum_{\lambda\in\mathbb{Z}_{\geq 0}^{n}}\frac{1}{\Delta(\lambda^{\prime})}\det[\mathrm{e}^{2i\pi\alpha_{r}\lambda_{s}^{\prime}}]\det[\mathrm{e}^{2i\pi\beta_{r}\lambda_{s}^{\prime}}]\det[\mathrm{e}^{-2i\pi\gamma_{r}\lambda_{s}^{\prime}}]$
@@
(3.27) $J[\gamma|\alpha,\beta]=\frac{1}{(2\pi)^{n(n-1)/2}}\sum_{\lambda\in\mathbb{Z}_{\geq 0}^{n}}\frac{1}{\Delta(\lambda^{\prime})}\det[\mathrm{e}^{2i\pi\alpha_{r}\lambda_{s}^{\prime}}]\det[\mathrm{e}^{2i\pi\beta_{r}\lambda_{s}^{\prime}}]\det[\mathrm{e}^{-2i\pi\gamma_{r}\lambda_{s}^{\prime}}].$
@@
3.2. Link with quantum cohomology of the Grassmannian
The goal of this section is to link the density (3.26) with structure constants of the quantum cohomology ring of Grassmannians $QH^{\bullet}(\mathbb{G}r)$. These constants are the Gromov-Witten invariants, which are related to characters via [26, Cor. 6.2]. We refer the reader to [22] and [27] for an introduction to the subject.
For $N\geq n$, the ring $QH^{\bullet}(\mathbb{G}r(n,N))$ has an additive basis $(g^{d}\otimes\sigma_{\lambda},d\geq 0,\lambda\in\mathbb{Z}_{N-n}^{n},(N-n\geq\lambda_{1}\geq\dots\geq\lambda_{n}\geq 0))$. We will denote $c_{\lambda,\mu}^{\nu,d}$ the structure constants of this ring so that
$\sigma_{\lambda}\cdot\sigma_{\mu}=\sum_{\nu,d\geq 0}c_{\lambda,\mu}^{\nu,d}q^{d}\otimes\sigma_{\nu^{\lor}}.$
where the sum is over partitions $(\nu,d)\in\mathbb{Z}_{N-n}^{n}$ such that $|\lambda|+|\mu|+|\nu|=n(N-n)+Nd$. The structure coefficients $c_{\lambda,\mu}^{\nu,d}$ are the degree $d$ Gromov-Witten invariants associated to the Schubert cycles $\sigma_{\lambda},\sigma_{\mu},\sigma_{\nu^{\lor}}$, see [26, Cor. 6.2]. The main result of this section is Theorem 3.8 below.
@@ 3.2. Link with quantum cohomology of the Grassmannian
Let $(\alpha,\beta,\gamma)\in\mathcal{H}^{3}$. For each $N\geq 1$, let $(\lambda_{N},\mu_{N},\nu_{N})$ be three partitions in $\mathbb{Z}_{N-n}^{n}$ such that $|\lambda_{N}|+|\mu_{N}|+|\nu_{N}|=n(N-n)+dN$ for some $d\in\mathbb{Z}_{\geq 0}$ and such that $\frac{1}{N}\lambda_{N}=\alpha+o(1)$, $\frac{1}{N}\mu_{N}=\beta+o(1)$ and $\frac{1}{N}\nu_{N}=\gamma+o(1)$ as $N\rightarrow+\infty$. Then,
@@
The rest of this section is devoted to the proof of Theorem 3.8. In subsection 3.2.1 we prove a determinantal formula for coefficients $c_{\lambda,\mu}^{\nu,d}$ along with some results on the quantities involved in the expression. In subsection 3.2.2 we prove Theorem 3.8 using Lemmas 3.16 and 3.15.
@@ 3.2.1. Determinantal expression for $c_{\lambda,\mu}^{\nu,d}$
Let $\lambda,\mu,\nu$ such that $|\lambda|+|\mu|+|\nu|=n(N-n)+Nd$. Then,
(3.29) $c_{\lambda,\mu}^{\nu,d}=\frac{1}{N^{n}}\sum_{I\in I_{n,N}}\frac{\det[\mathrm{e}^{\frac{2i\pi I_{r}(\lambda_{s}+(s-1))}{N}}]\det[\mathrm{e}^{\frac{2i\pi I_{r}(\mu_{s}+(s-1))}{N}}]\det[\mathrm{e}^{-\frac{2i\pi I_{r}((\nu^{\lor})_{s}+(s-1))}{N}}]}{\Delta(\xi^{I})}$
@@
Let $\xi=\exp{(2i\pi/N)}$ and for $I\in I_{n,N}$, set $\xi^{I}=(\xi^{I_{1}},\dots,\xi^{I_{n}})$. Let $S_{\lambda}(x_{1},\dots,x_{n})=\frac{\det[x_{r}^{(\lambda_{s}+(s-1))},1\leq r,s\leq n]}{\Delta(x)}$ be the Schur function corresponding to the partition $\lambda$. Using [26, Corollary 6.2]:
(3.30) $c_{\lambda,\mu}^{\nu,d}=\frac{1}{N^{n}}\sum_{I\in I_{n,N}}S_{\lambda}(\xi^{I})S_{\mu}(\xi^{I})S_{\nu}(\xi^{I})\frac{|\Delta(\xi^{I})|^{2}}{S_{(N-n)}(\xi^{I})}.$
Moreover, by [26, eq. (4.3)], one has
$\frac{S_{\nu}(\xi^{I})}{S_{(N-n)}(\xi^{I})}=\overline{S_{\nu^{\lor}}(\xi^{I})}$
@@
$c_{\lambda,\mu}^{\nu,d}$ $=\frac{1}{N^{n}}\sum_{I\in I_{n,m}}S_{\lambda}(\xi^{I})S_{\mu}(\xi^{I})S_{\nu^{\lor}}(\overline{\xi^{I}})|\Delta(\xi^{I})|^{2}$
(3.31) $=\frac{1}{N^{n}}\sum_{I\in I_{n,m}}\frac{\det[\mathrm{e}^{\frac{2i\pi I_{r}(\lambda_{s}+(s-1))}{N}}]\det[\mathrm{e}^{\frac{2i\pi I_{r}(\mu_{s}+(s-1))}{N}}]\det[\mathrm{e}^{-\frac{2i\pi I_{r}((\nu^{\lor})_{s}+(s-1))}{N}}]}{\Delta(\xi^{I})}$
@@
(3.32) $F(I,\lambda,\mu,\nu,N)=\frac{\det[\mathrm{e}^{\frac{2i\pi I_{r}(\lambda_{s}+(s-1))}{N}}]\det[\mathrm{e}^{\frac{2i\pi I_{r}(\mu_{s}+(s-1))}{N}}]\det[\mathrm{e}^{-\frac{2i\pi I_{r}((\nu^{\lor})_{s}+(s-1))}{N}}]}{\Delta(\xi^{I})}.$
@@
Let $I\in\left(\frac{1}{2}\mathbb{Z}\right)^{n}$ and $a\in\frac{1}{2}\mathbb{Z}$. We still assume that $|\lambda|+|\mu|+|\nu|=N(N-n)+Nd$ for some $d\in\mathbb{Z}_{\geq 0}$. Then,
@@
$\exp{\left(a\frac{2i\pi}{N}\left(|\lambda|+|\mu|-|\nu^{\lor}|+n(n-1)/2\right)\right)}=\exp{\left(a\frac{i\pi n(n-1)}{N}\right)}$
since $|\lambda|+|\mu|-|\nu^{\lor}|=0\ [N]$. The Vandermonde in the denominator of $F(I+a,\lambda,\mu,\nu,N)$ is
@@
From Lemma 3.10, we can shift $I$ by $\frac{n-1}{2}$. In the following, we will assume that $10\leq I_{n}<\dots<I_{1}\leq N-$ 1 and that the $I$’s are in $\mathbb{Z}$. Denote by $J_{n,N}$ the set
@@
The translation action $\Phi_{N}$ of $\mathbb{Z}$ on $J_{n,N}$ is given by
@@
where the tuple $I+(l,\dots,l)[N]$ consist of elements $i_{1}+l[N],\dots,i_{n}+l[N]$ sorted in the decreasing order.
@@
In order to give some properties of orbits of $\Phi_{m}$, let us recall the lexicographic order on $\mathbb{Z}_{\geq 0}^{n}$. For $I,J\in\mathbb{Z}_{\geq 0}^{n}$, let $r^{*}=\inf_{1\leq r\leq n}I_{r}\neq J_{r}$, with the convention that $r^{*}=0$ if $I=J$. We say that $I>J$ if $I_{r^{*}}>J_{r^{*}}$ and $I<J$ if $I_{r^{*}}<J_{r^{*}}$. This defines a total order on $\mathbb{Z}_{\geq 0}^{n}$.
Let $Orbits(N)$ denote the orbits of $\Phi_{N}$. For $\Omega_{N}\in Orbits(N)$ an orbit of $\Phi_{N}$, denote $\min(\Omega_{N})$ its minimal element with respect to the lexicographic order. Then, necessarily, $(\min(\Omega_{N}))_{n}=0$, otherwise $\Phi(-1,\min(\Omega_{N}))$ would be an element of $\Omega_{N}$ strictly inferior to $\min(\Omega_{N})$. For an ordered $n$-tuple $I$ of $\{0,\dots,N-1\}^{n}$, let $\Omega(I,N)$ denote its orbit under the action of $\Phi_{N}$.
@@
Let $G_{I,N}$ denote the stabilizer of $I$ under $\Phi_{N}$. Then, $N\mathbb{Z}\subset G_{I,N}$ so that $G_{I,N}=p_{N}\mathbb{Z}$ for some $p_{N}\geq 1$ such that $p_{N}|N$. Denote $d(N)=N-I_{1}\geq 1$. Then, for $p_{N}\mathbb{Z}$ to be the stabilizer of $I$, one must have $d(N)\leq p_{N}$ (recall that $I_{n}=0$). However, as $I_{1}$ is fixed, for $N>2I_{1}$, $\frac{N}{2}<d(N)\leq p_{N}$ which implies that $p_{N}=N$ since $p_{N}|N$ which gives an orbit of cardinal $N$.
Let us show that $I$ is minimal in its orbit $\Omega(I,N)$ when $N$ is large enough. Take $N$ such that $I_{1}\leq\frac{N}{2}$. The only points $J$ in the orbit of $I$ such that $J_{n}=0$ different from $I$ are
@@
$=\sum_{I_{1}>\dots>I_{n}=0}\mathrm{1}_{I=\min(\Omega(I,N))}|\Omega(I,m)|F(I,\lambda,\mu,\nu,N)$
@@
Let $N\geq n$ and consider $\Omega\in Orbits(N)$. Denote $I=\min(\Omega)$ its minimal element. Assume for the sake of contradiction that $I_{1}>N-\frac{N}{n}$. Divide the circle in $n$ pieces that we denote $P_{1},\dots P_{n}$ of length $\frac{N}{n}$ starting from $I_{n}=0$, that is, viewing the circle as $[0,1]$,
$P_{j}=\left[j-1\frac{N}{n},j\frac{N}{n}\right[,0\leq j\leq n-1.$
Since $I_{1}>N-\frac{N}{n}$ and $I_{n}=0$, $I_{1}$ and $I_{n}$ both belong to the last piece $P_{n}$. There are $n-2$ remaining elements $I_{2}>\dots>I_{n-1}$ to be placed inside the $n-1$ unused segments $P_{1},\dots,P_{n-1}$ and $P_{n}$. Thus, there exists $11\leq j\leq n-$ 1 such that $I\cap P_{j}=\emptyset$. Take the maximal such $j$ and consider $r=\max\{l\in[1,n]:I_{l}\geq j\frac{N}{n}\}$ the first element after $P_{j}$ starting from 00 counterclockwise on the circle. We claim that

### v2

@@ 3.1. A first density formula
$\mathrm{d}\mathbb{P}[\gamma|\alpha,\beta]=$ $\frac{\Delta(\mathrm{e}^{2i\pi\gamma})sf(n-1)}{(2\pi)^{n-1}n!\Delta(\mathrm{e}^{2i\pi\alpha})\Delta(\mathrm{e}^{2i\pi\beta})}\sum_{\lambda\in\mathbb{Z}_{\geq 0}^{n-1}}\frac{1}{\Delta(\lambda^{\prime})}\det[\mathrm{e}^{2i\pi\alpha_{r}\lambda_{s}^{\prime}}]\det[\mathrm{e}^{2i\pi\beta_{r}\lambda_{s}^{\prime}}]\det[\mathrm{e}^{-2i\pi\gamma_{r}\lambda_{s}^{\prime}}]$
@@
(3.25) $J[\gamma|\alpha,\beta]=\frac{1}{(2\pi)^{n(n-1)/2}}\sum_{\lambda\in\mathbb{Z}_{\geq 0}^{n-1}}\frac{1}{\Delta(\lambda^{\prime})}\det[\mathrm{e}^{2i\pi\alpha_{r}\lambda_{s}^{\prime}}]\det[\mathrm{e}^{2i\pi\beta_{r}\lambda_{s}^{\prime}}]\det[\mathrm{e}^{-2i\pi\gamma_{r}\lambda_{s}^{\prime}}].$
@@
3.2. Link with quantum cohomology of the Grassmannians
The goal of this section is to link the volume function (3.25) with structure constants of the quantum cohomology ring of Grassmannians $QH^{\bullet}(\mathbb{G}r)$ in the same way as the volume function in the coadjoint case is related to the classical cohomology ring of Grassmannians, see [8]. The structure constants in the unitary case are the Gromov-Witten invariants, which are related to characters via [32, Cor. 6.2]. We refer the reader to [26] and [33] for an introduction to the subject.
For $N\geq n$, denoted by $\mathbb{Z}_{N-n}^{n}$ the set of partition $\lambda\in\mathbb{Z}^{n}$ such that $N-n\geq\lambda_{1}\geq\dots\geq\lambda_{n}\geq 0$. Then, the ring $QH^{\bullet}(\mathbb{G}r(n,N))$ has an additive basis $(q^{d}\otimes\sigma_{\lambda},d\geq 0,\lambda\in\mathbb{Z}_{N-n}^{n})$. We will denote by $c_{\lambda,\mu}^{\nu,d}$ the structure constants of this ring so that
$\sigma_{\lambda}\cdot\sigma_{\mu}=\sum_{\nu,d\geq 0}c_{\lambda,\mu}^{\nu,d}q^{d}\otimes\sigma_{\nu}.$
where the sum is over pairs $(\nu,d)\in\mathbb{Z}_{N-n}^{n}\times\mathbb{N}$ such that $|\lambda|+|\mu|=|\nu|+Nd$. The structure coefficients $c_{\lambda,\mu}^{\nu,d}$ are the degree $d$ Gromov-Witten invariants associated to the Schubert cycles $\sigma_{\lambda},\sigma_{\mu},\sigma_{\nu^{\lor}}$, see [32, Cor. 6.2]. The main result of this section is Theorem 3.8 below.
@@ 3.2. Link with quantum cohomology of the Grassmannians
Let $(\alpha,\beta,\gamma)\in\mathcal{H}_{reg}^{3}$. For each $N\geq 1$, let $(\lambda_{N},\mu_{N},\nu_{N})$ be three partitions in $\mathbb{Z}_{N-n}^{n}$ such that $|\lambda_{N}|+|\mu_{N}|=|\nu_{N}|+dN$ for some $d\in\mathbb{Z}_{\geq 0}$ and such that $\frac{1}{N}\lambda_{N}=\alpha+o(1)$, $\frac{1}{N}\mu_{N}=\beta+o(1)$ and $\frac{1}{N}\nu_{N}=\gamma+o(1)$ as $N\rightarrow+\infty$. Then,
@@
The rest of this section is devoted to the proof of Theorem 3.8. In subsection 3.2.1 we prove a determinantal formula for the coefficients $c_{\lambda,\mu}^{\nu,d}$ along with some results on the quantities involved in the expression. In subsection 3.2.2 we prove Theorem 3.8 using Lemmas 3.16 and 3.15.
@@ 3.2.1. Determinantal expression for $c_{\lambda,\mu}^{\nu,d}$
Let $\lambda,\mu,\nu$ such that $|\lambda|+|\mu|=|\nu|+Nd$. Then,
(3.27) $c_{\lambda,\mu}^{\nu,d}=\frac{1}{N^{n}}\sum_{I\in I_{n,N}}\frac{\det[\mathrm{e}^{\frac{2i\pi I_{r}(\lambda_{s}+(s-1))}{N}}]\det[\mathrm{e}^{\frac{2i\pi I_{r}(\mu_{s}+(s-1))}{N}}]\det[\mathrm{e}^{-\frac{2i\pi I_{r}(\nu_{s}+(s-1))}{N}}]}{\Delta(\xi^{I})}$
@@
Let $\xi=\exp{(2i\pi/N)}$ and for $I\in I_{n,N}$, set $\xi^{I}=(\xi^{I_{1}},\dots,\xi^{I_{n}})$. Let $S_{\lambda}(x_{1},\dots,x_{n})=\frac{\det[x_{r}^{(\lambda_{s}+(s-1))},1\leq r,s\leq n]}{\Delta(x)}$ be the Schur function corresponding to the partition $\lambda$. Using [32, Corollary 6.2]:
(3.28) $c_{\lambda,\mu}^{\nu,d}=\frac{1}{N^{n}}\sum_{I\in I_{n,N}}S_{\lambda}(\xi^{I})S_{\mu}(\xi^{I})S_{\nu^{\lor}}(\xi^{I})\frac{|\Delta(\xi^{I})|^{2}}{S_{(N-n)}(\xi^{I})}.$
Moreover, by [32, eq. (4.3)], one has
$\frac{S_{\nu^{\lor}}(\xi^{I})}{S_{(N-n)}(\xi^{I})}=\overline{S_{\nu}(\xi^{I})}$
@@
$c_{\lambda,\mu}^{\nu,d}$ $=\frac{1}{N^{n}}\sum_{I\in I_{n,m}}S_{\lambda}(\xi^{I})S_{\mu}(\xi^{I})S_{\nu}(\overline{\xi^{I}})|\Delta(\xi^{I})|^{2}$
(3.29) $=\frac{1}{N^{n}}\sum_{I\in I_{n,m}}\frac{\det[\mathrm{e}^{\frac{2i\pi I_{r}(\lambda_{s}+(s-1))}{N}}]\det[\mathrm{e}^{\frac{2i\pi I_{r}(\mu_{s}+(s-1))}{N}}]\det[\mathrm{e}^{-\frac{2i\pi I_{r}(\nu_{s}+(s-1))}{N}}]}{\Delta(\xi^{I})}$
@@
(3.30) $F(I,\lambda,\mu,\nu,N)=\frac{\det[\mathrm{e}^{\frac{2i\pi I_{r}(\lambda_{s}+(s-1))}{N}}]\det[\mathrm{e}^{\frac{2i\pi I_{r}(\mu_{s}+(s-1))}{N}}]\det[\mathrm{e}^{-\frac{2i\pi I_{r}(\nu_{s}+(s-1))}{N}}]}{\Delta(\xi^{I})}.$
@@
Let $I\in\left(\frac{1}{2}\mathbb{Z}\right)^{n}$ and $a\in\frac{1}{2}\mathbb{Z}$. We still assume that $|\lambda|+|\mu|=|\nu|+Nd$ for some $d\in\mathbb{Z}_{\geq 0}$. Then,
@@
$\exp{\left(a\frac{2i\pi}{N}\left(|\lambda|+|\mu|-|\nu|+n(n-1)/2\right)\right)}=\exp{\left(a\frac{i\pi n(n-1)}{N}\right)}$
since $|\lambda|+|\mu|-|\nu|=0\ [N]$. The Vandermonde in the denominator of $F(I+a,\lambda,\mu,\nu,N)$ is
@@
From Lemma 3.10, we can shift $I$ by $\frac{n-1}{2}$. In the following, we will assume that $10\leq I_{n}<\dots<I_{1}\leq N-$ 1 and that the $I\in\mathbb{Z}^{n}$. Denote by $J_{n,N}$ the set
@@
The translation action of $\mathbb{Z}$ on $J_{n,N}$ is given by
@@
where the tuple $I+(l,\dots,l)[N]$ consists of the sequence of elements $i_{1}+l[N],\dots,i_{n}+l[N]$ sorted in the decreasing order.
@@
In order to give some properties of orbits of $\Phi_{N}$, let us recall the lexicographic order on $\mathbb{Z}_{\geq 0}^{n}$. For $I,J\in\mathbb{Z}_{\geq 0}^{n-1}$, let $r^{*}=\inf_{1\leq r\leq n}I_{r}\neq J_{r}$, with the convention that $r^{*}=0$ if $I=J$. We say that $I>J$ if $I_{r^{*}}>J_{r^{*}}$ and $I<J$ if $I_{r^{*}}<J_{r^{*}}$. This defines a total order on $\mathbb{Z}_{\geq 0}^{n}$ and by restriction on $J_{n,N}$.
Let $Orbits(N)$ denote the orbits of the action of $\Phi_{N}$ on $J_{n,N}$. For $\Omega_{N}\in Orbits(N)$, denote by $\min(\Omega_{N})$ its minimal element with respect to the lexicographic order. Then, necessarily, $(\min(\Omega_{N}))_{n}=0$, otherwise $\Phi(-1,\min(\Omega_{N}))$ would be an element of $\Omega_{N}$ strictly inferior to $\min(\Omega_{N})$. For an ordered $n$-tuple $I$ of $\{0,\dots,N-1\}^{n}$, let $\Omega(I,N)$ denote its orbit under the action of $\Phi_{N}$.
@@
Let $G_{I,N}$ denote the stabilizer of $I$ under $\Phi_{N}$. Then, $N\mathbb{Z}\subset G_{I,N}$ so that $G_{I,N}=p_{N}\mathbb{Z}$ for some $p_{N}\geq 1$ such that $p_{N}|N$. Set $d(N)=N-I_{1}\geq 1$. Then, for $p_{N}\mathbb{Z}$ to be the stabilizer of $I$, one must have $d(N)\leq p_{N}$ (recall that $I_{n}=0$), for otherwise $I_{1}+p_{N}\not\in\{I_{2},\ldots,I_{n}\}$. However, as $I_{1}$ is fixed, for $N>2I_{1}$, $\frac{N}{2}<d(N)\leq p_{N}$ which implies with $p_{N}|N$ that $p_{N}=N$. Hence, for such $N$, the corresponding orbit has cardinal $N$.
Let us show that $I$ is minimal in its orbit $\Omega(I,N)$ when $N$ is large enough. Take $N$ such that $I_{1}<\frac{N}{2}$. The only points $J$ in the orbit of $I$ such that $J_{n}=0$ and $J\not=I$ are
@@
$=\sum_{0\leq I_{n}<\dots<I_{1}\leq N-1}\mathrm{1}_{I=\min(\Omega(I,N))}|\Omega(I,m)|F(I,\lambda,\mu,\nu,N)$
@@
[…]
$P_{j}=\left]j-1\frac{N}{n},j\frac{N}{n}\right],1\leq j\leq n.$
Since $I_{1}>N-\frac{N}{n}$ and $I_{n}=0$, $I_{1}$ and $I_{n}$ both belong to the last interval $P_{n}$. There are $n-2$ remaining elements $I_{2}>\dots>I_{n-1}$ to be placed inside the $n-1$ unused intervals $P_{1},\dots,P_{n-1}$ and $P_{n}$. Thus, there exists $11\leq j\leq n-$ 1 such that $I\cap P_{j}=\emptyset$. Take the maximal such $j$ and consider $r=\max\{l\in[1,n]:I_{l}\geq j\frac{N}{n}\}$ the index of the smallest element of $I$ greater than $P_{j}$. We claim that